37. In a perfectly competitive and constant-cost industry demand is given by Q = 2000 - 2P. The short total cost function at the scale of production that minimizes long-run costs for each identical firm is TC = 1000 + 100q + 10q2.


A. What will the long-run price be in this industry? How many firms will there be in long-run equilibrium?


B. What is the equation of the short-run supply curve if the industry is in long-run equilibrium?


C. Suppose that demand in this industry increases to: Q = 4000 - 2P. What will the new long run price be in this industry and how many firms will there be?


d. Describe how the industry will adjust to reach the new long-run equilibrium and illustrate your answer by drawing a graph of the firm and a graph of the industry.



1
Expert's answer
2022-02-14T14:47:06-0500

a) ATC= MC=P

ATC="\frac{1000}{q}+ 100+10q"

MC="100+ 20q"

"\frac{1000}{q}+ 100+10q= 100+ 20q"


"\frac{1000}{q}+10q-20q= 0"

"10q^2= 1000"

"q^2= 100"

q= 10

P= 100+20(10)= 300

Number of firms in the industry

market Quantity (Q*) = 2000- 2(300)= 1400

nq= Q*

n="\frac{Q*}{q}=\frac {1400}{10}= 140"


b) minLRATC=LR supply


c) New Price

Q* 4000- 2(300)= 2400

n= "\frac{2400}{10}= 240"

d)